# m m OTr_ #1#3(#2#3)

### v2

@@ Abstract
What is fundamentally quantum? We argue that most of the features, problems, and paradoxes – such as the measurement problem, the Wigner’s friend paradox and its proposed solutions, single particle nonlocality, and no-cloning – allegedly attributed to quantum physics have a clear classical analogue if one is to interpret classical physics as fundamentally indeterministic. What really characterizes quantum physics boils down only to phenomena that involve $\hbar$, i.e., incompatible observables.
I Introduction
@@
In what is by now a most famous dictum, Richard Feynman stated that one “can safely say that nobody understands quantum mechanics”. This perhaps remains true even after a century since the inception of this theory. However, is this lack of understanding really unique to quantum theory, or were some fundamental issues already present, albeit somewhat hidden, in classical physics? Better said, are the fundamental features that characterize quantum physics – allegedly regarded as setting a watershed from the past of classical physics – truly unique to quantum theory? [^1] Questions along the same lines are addressed by previous works aimed at scaling down the fundamental difference between classical and quantum theory, such as [1, 2, 3, 4].
@@
In recent years, a number of works have challenged the view that fundamental indeterminacy was a novelty of the quantum world. On the contrary, classical physics can be regarded as a genuinely indeterministic theory [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 3, 20]. In particular, we have shown that the idea that classical physics is a deterministic theory is based on the tacit assumption that we called principle of infinite precision [15], namely, that there exists an actual value of every physical quantity, with its infinite determined information (represented by real numbers). If this strong assumption is instead relaxed, one introduces ontic indeterminacy [21, 22, 23], i.e., physical pure states contain at any point in time only finite information. In this way, when a measurement of higher precision is performed on a system, it generates genuinely new information that did not exist before. While indeterminacy is a kinematical property, at the dynamical level, even under classical symplectic evolution, this leads to indeterminism, i.e., the same state can generally evolve into different possible future outcomes [11, 18].
@@
This therefore shows that one of the supposedly main characteristic features of quantum physics, fundamental indeterminism, is not necessarily purely quantum. Here we build on this intuition, trying to identify as much as possible the genuine features of quantum physics and what can instead be achieved by an indeterministic interpretation of classical theory. Starting from the measurement problem (Sect. III), which is shared by a whole class of indeterministic theories, we show that there are also classical analogues of Wigner’s friend scenarios (Sect. III.1). Moreover, also this “classical measurement problem” can be solved by a number of interpretations that have exact parallels with the solutions provided by the prominent interpretations of quantum physics (Sect. IV).
@@
Additionally, in Sect. V, we discuss the possibility of finding single particle nonlocality à la Einstein and the classical analogue of entanglement which, however, does not lead to Bell nonlocality. [^2] These two types of nonlocality are related to what H. Wiseman distinguishes as two Bell theorems: the first states that “there exist quantum phenomena for which there is no theory satisfying locality and determinism”, whereas the second one rules out local causality [24]. We show that local causality cannot be ruled out in a classical scenario, even in the presence of ontic indeterminacy. Nonlocality à la Einstein concerns the problem of how a detector can “decide” to click while all others remain silent when measuring a single particle in a superposition of different locations; we show that this is achievable in classical indeterministic theories too. On the other hand, while we find a classical analogue of entanglement, this cannot lead to the violation of Bell inequalities due to the absence of incompatible variables. Finally, in Sect. VI, we show that the no-cloning theorem also holds in (indeterministic) classical physics.
@@
Thus, the conceptual gap between classical and quantum theories should be significantly reduced. However, quantum theory features incompatible observables that lead to effects unattainable by classical theory, even under an indeterministic interpretation thereof. In general terms, incompatibility refers to pairs of physical variables that cannot be simultaneously determined with arbitrary precision. When one physical variable becomes more defined (for instance, due to a measurement), this generally affects the (in)determinacy of all variables incompatible with it. In quantum mechanics, incompatibility is associated with non-commuting observables, a relationship encapsulated by the Heisenberg uncertainty principle: the product of their indeterminacies (i.e., standard deviations) is bounded by a fundamental lower limit specific to the observables in question, in proportion to Planck’s reduced constant $\hbar$. For example, in the case of position and momentum, this lower bound is /2Planck-constant-over-$\hbar/2$. The situation is drastically different in classical theory, even in an indeterministic interpretation thereof, where states in phase space exhibit ontic indeterminacy. In fact, there is no posited lower limit to the joint determinacy that variables can acquire (again due to, e.g., measurements). In particular, position and momentum in phase space can be increasingly determined independently of each other, up to arbitrary, though always finite, precision.
@@
We so identify the essence of quantum mechanics in phenomena involving Planck’s reduced constant, $\hbar$, that is, incompatible observables at the kinematical level. Heisenberg’s uncertainty relations, weak values [25], and the violation of Bell’s inequality, for example, rely on incompatible measurements interconnected by $\hbar$, and indeed lack classical analogues. It is of course very well-established that $\hbar$ characterizes quantum theory, but here we shed new light on the fact that, contrarily to the widespread belief, many features and paradoxes are not specific to quantum physics, unless they involve $\hbar$.
II Fundamental indeterminacy
@@
Figure 1: States of “orthodox” classical mechanics (left), where a physical state is a mathematical point, and of the indeterministic alternative interpretation of classical physics (right), based on finite information quantities (FIQs).
@@
Relaxing the principle of infinite precision leads to an alternative interpretation of classical physics which upholds fundamental indeterminacy. This means that pure states of classical physics are not mathematical points. We take this to be ontic indeterminacy, meaning that there exists no additional information in the universe that can specify the state of a system more precisely than what is provided by finite information (see Fig. 1). Even keeping unchanged the equations of motion, this (kinematic) indeterminacy leads to (dynamical) indeterminism: given a pure state, this has the potential to evolve into different future states, of which eventually only one will actualize. In this way, the standard or “orthodox” interpretation of classical physics in terms of dimensionless points (described by n-tuples of real numbers) is regarded as a hidden variable completion of indeterministic classical physics [12] (see Sect. IV).
The relinquishing of determinism in favor of an indeterministic description is usually considered the prerogative of quantum physics, which supposedly marks a drastic departure from classical physics. However, in what follows, we will review a specific model of classical indeterministic physics, based on what we called “finite information quantities” (FIQs) [11]. [^3] Note that our findings which we will discuss throughout this paper do not rely on our specific formalization of classical indeterminism, and can be equally derived from any model that upholds a deterministic dynamical evolution with stochastic actualizations. We will show how going down this path leads to recovering most of the issues believed to be characteristic of quantum physics, which now belong to a whole class of indeterministic theories.
II.1 Finite information quantities (FIQs)
In the orthodox interpretation of classical theory, a physical quantity $\gamma$ (say the position of a particle in one dimension) take value in the real numbers. Say that $\gamma\in[0,1]$ such that we can write it in binary base as follows:
@@ II.1 Finite information quantities (FIQs)
A propensity $q_{j}\in[0,1]\cap\mathbb{Q}$ monotonically quantifies the tendency of the $j$th bit to eventually take the value 1. This is to be interpreted as an objective, causally determined (but not by necessity) [18], intrinsic tendency for individual events to obtain. As notable examples, if $q_{j}=1$ (resp. 00) means that the $j$th bit will take value 1 (resp. 00) with certainty. If instead a bit has an associated propensity of 1/2, it means that the bit is totally indeterminate.
Propensities share most features of probabilities, but they differ insofar as they bear a causal structure, which at the formal level leads to substantial differences, like the rejection of standard Bayes’ rule. [^4] This distinguishes our propensities from the original proposal of Popper [26], for the latter conceived them as an interpretation of probabilities. We have discussed in detail this and other properties of propensities in Ref. [18].
@@
An example of a physical quantity $\gamma$ that fulfils the above definition of FIQs can thus be written as: [^5] Note that despite its intuitiveness and simplicity, this way of introducing FIQs is not fully satisfactory since the information content is not invariant under change of units and basic arithmetic [27]. A complete definition of FIQs is provided in the language of intutionism [28, 29, 16].
@@
In this way, at a given time $t$, the first $N$ bits are determined (i.e. the propensities are all either 0 or 1), making this FIQ indistinguishable from a standard real number up to this level of precision. However, the following bits are not yet actualized, i.e., they are still indeterminate. Among those, according to this simple model, a finite amount thereof (i.e., until the $M$-th bit) have a biased tendency to actualize their value to either 0 or 1, i.e., the propensities associated to each bits are $q_{k}\in(0,1)\cap\mathbb{Q}$, for every $\ N<k\leq M$. Finally, after the $M$-th bit, all the associated propensities are completely random, i.e., $q_{M+1},q_{M+2},\dots$ all equal to 1/2, ensuring that the condition of finite information is met (the information content I(1/2)=0). We stress that in this interpretation the pure state (i.e., the one containing maximal information) of a system is a list of potential properties. Hence, two systems in the same pure state will undergo, in general, different future evolutions depending on the actualization of the propensities as (creative) time passes [20].
III The measurement problem
Generally speaking, the quantum measurement problem can be formulated as the tension between the deterministic dynamical laws of a theory and the fact that we observe probabilistic (single) outcomes when we perform measurements. In fact, quantum mechanics is governed by the Schrödinger equation which uniquely maps states at different times via unitary transformations. However, the linearity of the Schrödinger equation allows for superpositions of different evolutions, while it is a very corroborated empirical fact that, in the ideal case, we find a single though probabilistic outcome corresponding to a single eigenvalue of the measurement operator every time we observe a system, i.e., we perform a measurement (whereas in general this corresponds to a POVM). Therefore, one postulates a “collapse” or state update, but the theory does not prescribe how and when this happens. Hence, the Schrödinger evolution is somehow effectively interrupted when the system interacts with a bunch of atoms which has been labeled “measurement apparatus”.
@@ III The measurement problem
The same situation one finds in alternative classical mechanics, where the state encapsulating the fundamental indeterminacy (via the collection of propensities in a FIQ) is evolved via Hamiltonian dynamics and uniquely mapped to complex branched configurations in phase space (for certain systems, such as chaotic systems), which correspond to arbitrarily large spread of the state in position and/or momentum under the constraint of Liouville’s theorem. Yet, every time we observe a system we find it localized, so there must be an actualization of the state that reduces the indeterminacy. As a simple example, consider a particle of approximate diameter $d$ (i.e., of physical size $d$ with some ontic indeterminacy) moving in one-dimension on a constrained segment $[0,\ell]$ (with walls that can be considered perfectly elastic). Assume that the initial velocity (say pointing to the left) has some ontic indeterminacy, i.e. $v_{0}\leq v(t=0)=v_{0}+\Delta v_{0}$. According to the laws of classical mechanics, the indeterminacy on the particle’s positions increases linearly as time passes, i.e. $\Delta x(t)=t\Delta v_{0}$. This means that there always exists a critical time $t_{c}:=\ell/\Delta v_{0}$, such that $\Delta x(t=t_{c})=\ell$. This means that, for any finite (even arbitrarily small) indeterminacy on the velocity, after finite time, the indeterminacy on the particle’s position saturates the whole space. However, if measured, one expects the particle to become localized in a region of a size $R\sim d\ll\ell$.
@@
While in quantum mechanics unitarity preserves the norm of vectors in the Hilbert space, in classical physics, symplectic dynamics preserves the volume in phase space (Liouville’s theorem). But in both theories, when there is a measurement or any analog of such a process, these standard dynamics are interrupted and we say that there is a “collapse” of the quantum state and, respectively, an actualization of some bits, hence reducing the indeterminacy of the classical indeterminate state.
@@
Therefore, at a more conceptual level, the measurement problem can be formulated as the questions: How does a single value of a physical variable become actualized out of its possible ones? (see Fig. 2). [^6] When we speak about a single value, we mean an effective, coarse-grained, approximate value, for also after actualization the fundamental indeterminacy would never completely disappear, i.e., there would never be determination with infinite information. How does potentiality become actuality? (see also [11, 17, 18]). It is well known that the answer to these questions is the main point of disagreement between different interpretations of quantum physics. Each interpretation provides a different “mechanism” that “explains” the observed state update, i.e., that “forces” the potentiality to become a determined outcome upon measurement.
III.1 Wigner’s friend

### v3

# Which features of quantum physics are not fundamentally quantum but are due to indeterminism?

@@ Abstract
What is fundamentally quantum? We argue that most of the features, problems, and paradoxes – such as the measurement problem, the Wigner’s friend paradox and its proposed solutions, single particle nonlocality, and no-cloning – allegedly attributed to quantum physics have a classical analogue if one is to interpret classical physics as fundamentally indeterministic. What really characterizes non-classical effects are incompatible physical quantities, which, in quantum quantum theory are associated to the fundamental constant $\hbar$.
1 Introduction
@@
In what is by now a most famous dictum, Richard Feynman stated that one “can safely say that nobody understands quantum mechanics". This perhaps remains true even after a century since the inception of this theory. However, is this lack of understanding really unique to quantum theory, or were some fundamental issues already present, albeit somewhat hidden, in classical physics? Better said, are the fundamental features that characterize quantum physics – allegedly regarded as setting a watershed from the past of classical physics – truly unique to quantum theory? [^1] Questions along the same lines are addressed by previous works aimed at scaling down the fundamental difference between classical and quantum theory, such as [1, 2, 3, 4, 5, 6].
@@
In recent years, a number of works have challenged the view that fundamental indeterminacy was a novelty of the quantum world. On the contrary, classical physics can be regarded as a genuinely indeterministic theory [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 4, 23]. In particular, we have shown that the idea that classical physics is a deterministic theory is based on the tacit assumption that we called principle of infinite precision [18], namely, that there exists an actual value of every physical quantity, with its infinite determined information (represented by real numbers). If this strong assumption is instead relaxed, one introduces ontic indeterminacy [24, 25, 26], i.e., the description in the theory of physical pure states contain at any time only finite information. In this way, when a measurement of higher precision is performed on a system, it generates genuinely new information that did not exist before. While indeterminacy is a kinematical property, at the dynamical level, even under classical symplectic evolution, this leads to indeterminism, i.e., the same state can generally evolve into different possible future outcomes [13, 21].
@@
This therefore shows that one of the supposedly main characteristic features of quantum physics, fundamental indeterminism, is not necessarily purely quantum. Here we build on this intuition, trying to identify as much as possible the genuine features of quantum physics and what can instead be achieved by an indeterministic interpretation of classical theory. Starting from the measurement problem (Sect. 3), which is shared by a whole class of indeterministic theories, we show that there are also classical analogues of Wigner’s friend scenarios (Sect. 3.1). Moreover, also this “classical measurement problem" can be solved by a number of interpretations that have exact parallels with the solutions provided by the prominent interpretations of quantum physics (Sect. 4).
@@
Additionally, in Sect. 5, we discuss the possibility of finding single particle nonlocality à la Einstein and the classical analogue of entanglement which, however, does not lead to Bell nonlocality. [^2] These two types of nonlocality are related to what H. Wiseman distinguishes as two Bell theorems: the first states that “there exist quantum phenomena for which there is no theory satisfying locality and determinism”, whereas the second one rules out local causality [27]. We show that local causality cannot be ruled out in a classical scenario, even in the presence of ontic indeterminacy. Nonlocality à la Einstein concerns the problem of how a detector can “decide" to click while all others remain silent when measuring a single particle in a superposition of different locations; we show that this is achievable in classical indeterministic theories too. On the other hand, while we find a classical analogue of entanglement, this cannot lead to the violation of Bell inequalities due to the absence of incompatible physical quantities. Finally, in Sect. 6, we show that the no-cloning theorem also holds in (indeterministic) classical physics.
@@
Thus, the conceptual gap between classical and quantum theories is significantly reduced. However, quantum theory features incompatible physical quantities (also called variables) that lead to effects unattainable by classical theory, even under an indeterministic interpretation thereof. In general terms, incompatibility refers to pairs of physical quantities that cannot be simultaneously determined with arbitrary precision (at the ontological level). When one physical variable becomes more defined (for instance, due to a measurement), this generally affects the (in)determinacy of all variables incompatible with it. This can actually serve as a definition: a theory is considered classical if it does not feature incompatible physical quantities at the ontological level.
Quantum mechanics, on the contrary, features incompatible quantities, and this is associated in the theory with non-commuting observables. Such a relationship is encapsulated by the Heisenberg uncertainty principle: the product of their indeterminacies (i.e., standard deviations) is bounded by a fundamental lower limit specific to the observables in question, in proportion to Planck’s reduced constant $\hbar$. For example, in the case of position and momentum, this lower bound is /2Planck-constant-over-$\hbar/2$. We would like to stress that what makes quantum theory non-classical, from this point of view, is the fact that $\hbar>0$, while its particular numerical value – which, however, is characteristic of quantum theory – does not play a special role for the incompatibility of observables. The situation is drastically different in classical theory, even in an indeterministic interpretation thereof, where states in phase space exhibit ontic indeterminacy. In fact, there is no posited lower limit to the joint determinacy that variables can acquire (again due to, e.g., measurements). In particular, position and momentum in phase space can be increasingly determined independently of each other, up to arbitrary, though always finite, precision.
@@
We so identify the essence of quantum mechanics in phenomena involving Planck’s reduced constant, $\hbar$, that is, a strictly positive constant that guarantees the existence of incompatible quantities at the kinematical level. Heisenberg’s uncertainty relations, weak values [28], and the violation of Bell’s inequality, for example, rely on incompatible measurements interconnected by $\hbar$, and indeed lack classical analogues. It is of course very well-established that $\hbar$ characterizes quantum theory, but here we shed new light on the fact that, contrarily to the widespread belief, many features and paradoxes are not specific to quantum physics, unless they involve incompatible physical quantities, which in quantum theory are related to each other by $\hbar$.
2 Fundamental indeterminacy
@@
Figure 1: States of “orthodox" classical mechanics (left), where a physical state is a point in phase space, and of the indeterministic alternative interpretation of classical physics (right), based on finite information quantities (FIQs).
@@
Relaxing the principle of infinite precision leads to an alternative interpretation of classical physics which upholds fundamental indeterminacy. This means that pure states of classical physics are not points in phase space. We take this to be ontic indeterminacy, meaning that there exists no additional information in the universe that can specify the state of a system more precisely than what is provided by finite information (see Fig. 1). Even keeping unchanged the equations of motion, this (kinematic) indeterminacy leads to (dynamical) indeterminism: given a pure state, this has the potential to evolve into different future states, of which eventually only one will actualize. In this way, the standard or “orthodox" interpretation of classical physics in terms of dimensionless points (described by n-tuples of real numbers) is regarded as a hidden variable completion of indeterministic classical physics [14] (see Sect. 4).
The relinquishing of determinism in favor of an indeterministic description is usually considered the prerogative of quantum physics, which supposedly marks a drastic departure from classical physics. However, in what follows, we will review a specific model of classical indeterministic physics, based on what we called “finite information quantities" (FIQs) [13]. [^3] Note that our findings which we will discuss throughout this paper do not rely on our specific formalization of classical indeterminism, and can be equally derived from any model that upholds a deterministic dynamical evolution with stochastic actualizations. We will show how going down this path leads to recovering most of the issues believed to be characteristic of quantum physics, which now belong to a whole class of indeterministic theories.
2.1 Finite information quantities (FIQs)
In the orthodox interpretation of classical theory, a physical quantity $\gamma$ (say the position of a particle in one dimension) takes value in the real numbers. Say that $\gamma\in[0,1]$ such that we can write it in binary base as follows:
@@ 2.1 Finite information quantities (FIQs)
A propensity $q_{j}\in[0,1]\cap\mathbb{Q}$ monotonically quantifies the tendency of the $j$th bit to eventually take the value 1. This is to be interpreted as objective, intrinsic tendencies for some events to occur [21]. As notable examples, if $q_{j}=1$ (resp. 00) means that the $j$th bit will take value 1 (resp. 00) with certainty. If instead a bit has an associated propensity of 1/2, it means that the bit is totally indeterminate.
Propensities share most features of probabilities, but they differ insofar as they bear a causal structure, which at the formal level leads to substantial differences. [^4] This distinguishes our propensities from the original proposal of Popper [29], for the latter conceived them as an interpretation of probabilities. We have discussed in detail this and other properties of propensities in Ref. [21].
@@
An example of a physical quantity $\gamma$ that fulfils the above definition of FIQs can thus be written as:
@@
In this way, at a given time $t$, the first $N$ bits are determined (i.e. the propensities are all either 0 or 1), making this FIQ indistinguishable from a standard real number up to this level of precision. However, the following bits are not yet actualized, i.e., they are still indeterminate. Among those, according to this simple model, a finite amount thereof (i.e., until the $M$-th bit) have a biased tendency to actualize their value to either 0 or 1, i.e., the propensities associated to each bits are $q_{k}\in(0,1)\cap\mathbb{Q}$, for every $\ N<k\leq M$. Finally, after the $M$-th bit, all the associated propensities are completely random, i.e., $q_{M+1},q_{M+2},\dots$ all equal to 1/2, ensuring that the condition of finite information is met (the information content I(1/2)=0). We stress that in this interpretation the pure state (i.e., the one containing maximal information) of a system is a list of propensities. To stress this difference, the pure state $(2/3,1/2,1/2,\dots)$ is, for example, different from a mixture of the two pure states $(1,1/2,1/2,\dots)$ and $(0,1/2,1/2,\dots)$ with probabilities $2/3$ and $1/3$, respectively. Two systems in the same pure state will undergo, in general, different future evolutions depending on the actualization of the propensities as (creative) time passes [23].
3 The measurement problem
Generally speaking, the quantum measurement problem can be formulated as the tension between the deterministic dynamical laws of a theory and the fact that we observe probabilistic (single) outcomes when we perform measurements. In fact, quantum mechanics is governed by the Schrödinger equation which uniquely maps states at different times via unitary transformations. However, the linearity of the Schrödinger equation entails that linear combinations of quantum states evolve coherently into other superposition states, while it is a very corroborated empirical fact that, in the ideal case, we find a single though probabilistic outcome corresponding to a single eigenvalue of the measurement operator every time we observe a system, i.e., we perform a measurement (whereas in general this corresponds to a POVM). Therefore, one postulates a “collapse" or state update, but the theory does not prescribe how and when this happens. Hence, the Schrödinger evolution is somehow effectively interrupted when the system interacts with a bunch of atoms which has been labeled “measurement apparatus".
@@ 3 The measurement problem
The same situation one finds in alternative classical mechanics, where the state encapsulating the fundamental indeterminacy (via the collection of propensities in a FIQ) is evolved via Hamiltonian dynamics and uniquely mapped to complex branched configurations in phase space (for certain systems, such as chaotic systems), which correspond to arbitrarily large spread of the state in position and/or momentum under the constraint of Liouville’s theorem. Yet, every time we observe a system we find it localized, so there must be an actualization of the state that reduces the indeterminacy. As a simple example, consider a particle of approximate diameter $d$ (i.e., of physical size $d$ with some ontic indeterminacy) moving in one-dimension on a constrained segment $[0,\ell]$ (with walls that can be considered perfectly elastic). Assume that the initial velocity (say pointing to the left) has some ontic indeterminacy, i.e. its state is the interval $v_{0}\pm\delta v_{0}$. This means that the propensities are all determined (i.e., have values either 0 or 1) up to the precision of $\delta v_{0}$, which quantifies the ontic indeterminacy. According to the laws of classical mechanics, the indeterminacy on the particle’s positions increases linearly as time passes, i.e. $\Delta x(t)=t\delta v_{0}$. This means that there always exists a critical time $t_{c}:=\ell/\delta v_{0}$, such that $\delta x(t=t_{c})=\ell$. This means that, for any finite (even arbitrarily small) indeterminacy on the velocity, after finite time, the indeterminacy on the particle’s position saturates the whole space. However, if measured, one expects the particle to become localized in a region of a size $R\sim d\ll\ell$.
@@
While in quantum mechanics unitarity preserves the norm of vectors in the Hilbert space, in classical physics, symplectic dynamics preserves the volume in phase space (Liouville’s theorem). But in both theories, when there is a measurement or any analog of such a process, these standard dynamics are interrupted and we say that there is a “collapse" of the quantum state and, respectively, an actualization of some bits, hence reducing the indeterminacy of the classical indeterminate state.
@@
Therefore, at a more conceptual level, the measurement problem can be formulated as the questions: How does a single value of a physical variable become actualized out of its possible ones? (see Fig. 2). [^5] When we speak about a single value, we mean an effective, coarse-grained, approximate value, for also after actualization the fundamental indeterminacy would never completely disappear, i.e., there would never be determination with infinite information. How does potentiality become actuality? (see also [13, 20, 21]). It is well known that the answer to these questions is the main point of disagreement between different interpretations of quantum physics. Each interpretation provides a different “mechanism” that “explains" the observed state update, i.e., that “forces” the potentiality to become a determined outcome upon measurement.
3.1 Wigner’s friend